OpenStudy (anonymous):

HELPPPPPP PLEASE !!!!!! simplify the polynomial 3x2+8x+7x

OpenStudy (henriquerb):

3x²+8x+7x = 3x²+15x = x(3x+15)

OpenStudy (henriquerb):

?? is that it?

OpenStudy (anonymous):

yes thank you

OpenStudy (henriquerb):

You are welcome